Sarah Aiken is the owner of Macaron Paws, a dog-grooming service. At standard, it takes 1.0 hour to groom each dog. During the month of October, it took Macaron Paws employees 368 hours to groom 282 dogs, at a total labor cost of $4,894. In November, employees spent 326 hours grooming 286 dogs, at a total labor cost of $5,901; in December, 335 hours were used to groom 349 dogs, at a total labor cost of $6,499. The wage rate standard for groomers is $13 per hour.

(a) Calculate the labor rate and efficiency variances for each of the past three months. (Round per unit amounts to 2 decimal places and final answers to 0 decimal places, e.g. 125. Enter all variances as a positive number.)

October November December
Labor Rate Variance $

Neither Unfavorable nor FavorableFavorableUnfavorable

$

UnfavorableNeither Unfavorable nor FavorableFavorable

$

UnfavorableFavorableNeither Unfavorable nor Favorable

Labor Efficiency Variance $

UnfavorableNeither Unfavorable nor FavorableFavorable

$

UnfavorableNeither Unfavorable nor FavorableFavorable

$

UnfavorableFavorableNeither Unfavorable nor Favorable
Answer #1

(a) Calculate the labor rate and efficiency variances for each of the past three months. (Round per unit amounts to 2 decimal places and final answers to 0 decimal places, e.g. 125. Enter all variances as a positive number.)

October November December
Labor Rate Variance (13*368-4894) = 110

Unfavorable

(13*326-5901) = 1663 Unfavorable (13*335-6499) = 2144 Unfavorable
Labor Efficiency Variance (282-368)*13 = 1118 Unfavorable (286-326)*13 = 520 Unfavorable (349-335)*13 = 182 Favorable